OpenStudy (anonymous):

which of the following is equivalent to the expression (t^2-T-6)(T-3)^-1 a T^3-9t+18 b (T-2)(t+3) c (T+2)(T-3) d(T-2) e (T+2) Anwser is E but dont know how to get it

OpenStudy (rane):

use your calculator and simplify all of them then c which has the same answer

OpenStudy (anonymous):

i dont know how to type it on the calc

OpenStudy (rane):

oh!

OpenStudy (rane):

its E it gives u 2 same as the expression in the q'

OpenStudy (anonymous):

oh whoops your right read the anwser sheet wrong

OpenStudy (rane):

ok

OpenStudy (rane):

do u have more q's?

OpenStudy (rane):

ok gd luck with the rest :)

OpenStudy (phi):

you have to know 2 things first \[(T-3)^{-1}= \frac{1}{(T-3)} \] second, you need to be able to factor \[ (T^2 -T -6) \] to factor, first look at the last number: -6 write a pairs that give you 6 when you multiply: 1,6 2,3 the minus sign on the -6 means, look at the difference of each pair 1-6 = -5 2-3 = -1 if you find a difference that matches the middle number -T (which is -1 T) use that pair: 2 and -3 write the factors (T+2)(T-3) now you have \[ \frac{(T+2)(T-3)}{(T-3)} = (T+2) \] because (T-3) divided by (T-3) is 1